OpenStudy (mattyboyy):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? A) (x + 2)^2 + (y + 3)^2 = 9 B) (x − 2)^2 + (y − 3)^2 = 16 C) (x − 2)^2 + (y − 3)^2 = 4 D) (x + 2)^2 + (y + 3)^2 = 16 E) (x − 2)^2 + (y + 3)^2 = 9

OpenStudy (mattyboyy):

@iGreen @jswagger @neonumbrella5115

OpenStudy (anonymous):

if the center is (p, q) and radius r then the standard form is\[(x-p)^2+(y-q)^2 = r^2\]

OpenStudy (anonymous):

(0+5)^2 + (0 - 1)^2 =? 26 25 + 1 =? 26 26 = 26. So it checks.

OpenStudy (anonymous):

for the center (-2,-3) and point (-2,0) on the circle the radius is simply the distance between these two. Using distance formula we'll get r = 3 so the equation of the circle is \[(x-(-2))^2 +(y-(-3))^2= 3^2\]\[(x+2)^2 +(y+3)^2 = 9\]

OpenStudy (anonymous):

so option A)